the position of a particle along a straight line is given by s=t^3-9t^2-15t here t is in seconds determine its maximum accleration during the time 0<=t<=10s

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear akash

s=t3-9t2-15t

v=ds/dt =3t2-18t-15

a=dv/dt =6t-18

maximum acelaation occure at t=10 sec

amax=6*10-18

        =42


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin